### v1

@@ 1 Introduction
In the context of the NCP, the set $\partial\Omega_{{\displaystyle\varepsilon}}=\cup_{j=1}^{N}\Omega_{{\displaystyle\varepsilon}_{j}}$ is a collection of NN small interior traps of extent $\mathcal{O}({\displaystyle\varepsilon})$ centered at $\textbf{x}_{j}$, such that $\Omega_{{\displaystyle\varepsilon}_{j}}\to\textbf{x}_{j}$ as ${\displaystyle\varepsilon}\to 0$ for $k=1,\ldots,N$. The boundary condition (4b) encodes a range of reactivities $0<\kappa_{j}<\infty$ for the absorber. When $\kappa_{j}\to\infty$, the trap becomes absorbing so that any particle reaching $\partial\Omega_{{\displaystyle\varepsilon}_{j}}$ reacts rapidly. Conversely for $\kappa_{j}\ll 1$, the trap is weakly reactive so that particles reaching $\partial\Omega_{{\displaystyle\varepsilon}_{j}}$ are mostly reflected [BressloffSchumm2022, PCB2025, CengizLawley2024, GrebenkovWard2025a]. In the case of the NEP, the capture set is a collection of absorbing or partially absorbing boundary windows (see Fig. 1).
@@ 2.1 Two Dimensional analysis
(16) $\tilde{\tau}_{0}=\frac{|\Omega|}{2\pi D\nu N}\left[1+\frac{2\pi}{N\nu}p(\textbf{x}_{1},\ldots,\textbf{x}_{N})\right]\qquad p(\textbf{x}_{1},\ldots,\textbf{x}_{N})=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathcal{G}_{i,j}.$
@@ 2.2 The Neumann eigenvalue problem for the Laplacian
In the limit as ${\displaystyle\varepsilon}\to 0$ for NN identical traps, the limiting behavior (cf. [Sarafa2021, KTW_2005]) of $\lambda_{0}$ is
@@
In two dimensions we discretize the integral equation (32) using a standard patch-based collocation method, in which the unknowns of the system are values of the unknown density at a set of discretization nodes. For ease of exposition, we write the integral equation (32) as $\frac{1}{2}\sigma+\mathcal{K}[\sigma]=f,$ where $\mathcal{K}_{\textbf{y}}$ is a compact weakly-singular integral operator. Let $K_{\textbf{y}}$ denote its kernel. In two dimensions, the boundary $\partial\Omega$ is divided into panels, each of which is discretized using a $k^{th}$ order Gauss-Legendre rule. Let $\{\textbf{x}_{j}\}_{j=1}^{N}$ denote the collection of discretization nodes on $\partial\Omega.$ We then enforce the integral equation at the discretization nodes, yielding a set of NN equations
@@
In order to fully discretize the system we require suitable quadrature nodes $w_{j,k}$ for computing the integrals on the left-hand side of (36). For points which are far away, we can choose $w_{j,k}$ independent of $j$. Typically these quadrature weights $w_{k}$ are chosen so that the quadrature rule with nodes $\textbf{x}_{j}$ and weights $w_{j}$ accurately integrate sufficiently smooth functions on $\partial\Omega.$ For $\textbf{x}_{j}$ near $\textbf{x}_{k},$ custom quadrature formulae based on a mixture of adaptive integration and generalized Gaussian quadratures are used. Let $\mathcal{S}_{j}$ denote the set of $k$ such that $w_{j,k}$ requires such a correction. We note that if patches are approximately uniform in size, that $\mathcal{S}_{j}$ will typically be $\mathcal{O}(1)$ for large NN. Let $W_{j,k}$ denote the corresponding corrections. We can then replace (36) by the approximation
@@ 4 Results
In [Sarafa2021], the explicit solution of (1) in elliptical domains was used to identify a bifurcation of optimal placement away from linear alignment. Specifically, the optimal arrangement of traps corresponds to equally spaced placement along the semi-major axis at small NN before bifurcation off the semi-major axis at a critical eccentricity. In Fig. 7, we plot results for $N=4$ traps and a ellipse of area $|\Omega|=\pi ab=\pi$ for various semi-major axis $a$ and $b=1/a$.
@@
Our process is to sample standard normal random variables for $\{a_{k}\}_{k=1}^{M}$ and $\{b_{k}\}_{k=1}^{M}$ then set $a_{0}=1.1\times\sum_{k=1}^{N}\big(|a_{k}|+|b_{k}|\big)$. In Fig. 8, we show the result of numerical optimization of (7) for trap numbers $N=1:20$. At smaller NN we observe a tendency of traps to be located near the center and the geometric lobes of the region. For large numbers, we see that the traps become almost equally spaced throughout the domain.
@@
As the number NN of boundary traps increases, the optimization problem becomes influenced by interaction between traps and in particular their pairwise distance through the logarithmic free space Green’s function. Interestingly, for $N=2$ we see in the ellipse case $a=2$ that the optimal locations are not the points $(0,\pm b)$ as would be suggested by a preference towards points of lowest boundary curvature. Instead, the optimal configuration moves away from regions of preferred curvature in order to increase the pairwise distance. However, we do observe in this elliptical example that points have a preference to cluster around these two points of most negative curvature.
@@ 5 Discussion
In future work, we will also address numerical algorithms for the corresponding three dimensional Green’s function. In the case where NN small traps are located inside a bounded three dimensional region $\Omega$, similar asymptotic analysis have been performed in the limit as ${\displaystyle\varepsilon}\to 0$ [CHEVIAKOV2011, BL2018]. In the case that each trap has a different size or shape, the GMFPT is given (see Principal Result 3.1 of [CHEVIAKOV2011]) by

### v2

@@ 1 Introduction
In the context of the NCP, the set $\partial\Omega_{{\displaystyle\varepsilon}}=\cup_{j=1}^{N}\Omega_{{\displaystyle\varepsilon}_{j}}$ is a collection of $N$ small interior traps of extent $\mathcal{O}({\displaystyle\varepsilon})$ centered at $\textbf{x}_{j}$, such that $\Omega_{{\displaystyle\varepsilon}_{j}}\to\textbf{x}_{j}$ as ${\displaystyle\varepsilon}\to 0$ for $k=1,\ldots,N$. The boundary condition (4b) encodes a range of reactivities $0<\kappa_{j}<\infty$ for the absorber. When $\kappa_{j}\to\infty$, the trap becomes absorbing so that any particle reaching $\partial\Omega_{{\displaystyle\varepsilon}_{j}}$ reacts rapidly. Conversely for $\kappa_{j}\ll 1$, the trap is weakly reactive so that particles reaching $\partial\Omega_{{\displaystyle\varepsilon}_{j}}$ are mostly reflected [BressloffSchumm2022, PCB2025, CengizLawley2024, GrebenkovWard2025a]. In the case of the NEP, the capture set is a collection of absorbing or partially absorbing boundary windows (see Fig. 1).
@@ 2.1 Two Dimensional analysis
(16) $\tilde{\tau}_{0}=\frac{|\Omega|}{2\pi D\nu N}\left[1+\frac{2\pi\nu}{N}p(\textbf{x}_{1},\ldots,\textbf{x}_{N})\right]\qquad p(\textbf{x}_{1},\ldots,\textbf{x}_{N})=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathcal{G}_{i,j}.$
@@ 2.2 The Neumann eigenvalue problem for the Laplacian
In the limit as ${\displaystyle\varepsilon}\to 0$ for $N$ identical traps, the limiting behavior (cf. [Sarafa2021, KTW_2005]) of $\lambda_{0}$ is
@@
In two dimensions we discretize the integral equation (32) using a standard patch-based collocation method, in which the unknowns of the system are values of the unknown density at a set of discretization nodes. For ease of exposition, we write the integral equation (32) as $\frac{1}{2}\sigma+\mathcal{K}[\sigma]=f,$ where $\mathcal{K}_{\textbf{y}}$ is a compact weakly-singular integral operator. Let $K_{\textbf{y}}$ denote its kernel. In two dimensions, the boundary $\partial\Omega$ is divided into panels, each of which is discretized using a $k^{th}$ order Gauss-Legendre rule. Let $\{\textbf{x}_{j}\}_{j=1}^{N}$ denote the collection of discretization nodes on $\partial\Omega.$ We then enforce the integral equation at the discretization nodes, yielding a set of $N$ equations
@@
In order to fully discretize the system we require suitable quadrature nodes $w_{j,k}$ for computing the integrals on the left-hand side of (36). For points which are far away, we can choose $w_{j,k}$ independent of $j$. Typically these quadrature weights $w_{k}$ are chosen so that the quadrature rule with nodes $\textbf{x}_{j}$ and weights $w_{j}$ accurately integrate sufficiently smooth functions on $\partial\Omega.$ For $\textbf{x}_{j}$ near $\textbf{x}_{k},$ custom quadrature formulae based on a mixture of adaptive integration and generalized Gaussian quadratures are used. Let $\mathcal{S}_{j}$ denote the set of $k$ such that $w_{j,k}$ requires such a correction. We note that if patches are approximately uniform in size, that $\mathcal{S}_{j}$ will typically be $\mathcal{O}(1)$ for large $N$. Let $W_{j,k}$ denote the corresponding corrections. We can then replace (36) by the approximation
@@ 4 Results
In [Sarafa2021], the explicit solution of (1) in elliptical domains was used to identify a bifurcation of optimal placement away from linear alignment. Specifically, the optimal arrangement of traps corresponds to equally spaced placement along the semi-major axis at small $N$ before bifurcation off the semi-major axis at a critical eccentricity. In Fig. 7, we plot results for $N=4$ traps and a ellipse of area $|\Omega|=\pi ab=\pi$ for various semi-major axis $a$ and $b=1/a$.
@@
Our process is to sample standard normal random variables for $\{a_{k}\}_{k=1}^{M}$ and $\{b_{k}\}_{k=1}^{M}$ then set $a_{0}=1.1\times\sum_{k=1}^{N}\big(|a_{k}|+|b_{k}|\big)$. In Fig. 8, we show the result of numerical optimization of (7) for trap numbers $N=1:20$. At smaller $N$ we observe a tendency of traps to be located near the center and the geometric lobes of the region. For large numbers, we see that the traps become almost equally spaced throughout the domain.
@@
As the number $N$ of boundary traps increases, the optimization problem becomes influenced by interaction between traps and in particular their pairwise distance through the logarithmic free space Green’s function. Interestingly, for $N=2$ we see in the ellipse case $a=2$ that the optimal locations are not the points $(0,\pm b)$ as would be suggested by a preference towards points of lowest boundary curvature. Instead, the optimal configuration moves away from regions of preferred curvature in order to increase the pairwise distance. However, we do observe in this elliptical example that points have a preference to cluster around these two points of most negative curvature.
@@ 5 Discussion
In future work, we will also address numerical algorithms for the corresponding three dimensional Green’s function. In the case where $N$ small traps are located inside a bounded three dimensional region $\Omega$, similar asymptotic analysis have been performed in the limit as ${\displaystyle\varepsilon}\to 0$ [CHEVIAKOV2011, BL2018]. In the case that each trap has a different size or shape, the GMFPT is given (see Principal Result 3.1 of [CHEVIAKOV2011]) by